The following pension-related data pertain to Metro Recreation's noncontributory, defined benefit pension plan for 2018: ($ in 000s) Jan. 1 Dec. 31 Projected benefit obligation $ 4,800 $ 5,080 Accumulated benefit obligation 3,750 4,020 Plan assets (fair value) 5,580 6,025 Interest (discount) rate, 7% Expected return on plan assets, 10% Prior service cost?AOCI (from Dec. 31, 2017, amendment) 910 Net loss?AOCI 578 Average remaining service life: 10 years Gain due to changes in actuarial assumptions 52 Contributions to pension fund (end of year) 410 Pension benefits paid (end of year) 365 Required: Prepare a pension spreadsheet that shows the relationships among the various pension balances, shows the changes in those balances, and computes pension expense for 2018. (Enter credit amounts with a minus sign and debit amounts with a positive sign. Enter your answers in thousands (i.e. 200,000 should be entered as 200).)
Answer #1

Answer:

The elements of pension plan are projected benefit obligation, plan asset, prior service cost, gains and losses, periodic pension expense and fund status of the plan. The projected benefit obligation includes accrued interest, revised estimates, plan amendments, and payment of benefits. The plan assets include charges due to investment returns, employer contributions, and the payment of benefits. These elements of the pension plan are interrelated among each other. These can be brought together usin Pension Spreadsheet.

Pension Spreadsheet:

Particulars($in,000s)

PBO

Plan Assets

Prior service cost-AOCI

Net Loss-AOCI

Pension Expense

Cash

Net Pension (liability)/Asset

Balance, Jan.1,2018

(4100)

4530

840

477

430

Service Cost

(332)

332

(332)

Interest Cost, 7%

(287)

287

(287)

Expected Return on Assets

453

(453)

453

Adjust for: Loss on assets

(53)

53

(53)

Amortization of:

Prior service cost

(70)

70

Net loss

(2)

2

Gain o PBO

44

(44)

44

Cash Funding

340

(340)

340

Retire Benefits

295

(295)

Balance Dec31,2018

(4380)

4975

770

484

238

595

Workings:

  1. Interest Cost=Rate X Projected Benefit Obligation

                        =7% X 4100=287

  1. Service cost:

Particulars

Amount($ in 000s)

PBO, Dec 31

4380

Less: PBO, Jan 1

4100

Less: Interest Cost

287

Gain due to changes in actuarial assumption

44

Pension benefits paid ( end of the year)

295

Service Cost

332

3. Expected Rate on Assets=Plan assets X return on assets

= 4530 X 10%=453

4.Loss on Assets= Expected return on assets- Actual return on assets

=453-400=53

5.Prior Service Cost=(Prior Service Cost-AOCI)/Service Years

=840/12=70

6.Net Loss=(Net Loss-AOCI-Expected return on assets)/Service Life

=(477-453)/12=2

Abbreviations:

PBO- Projected Benefit Obligation

AOCI- Accumulated Other Comprehensive Income

OCI-Other Comprehensive Income